OpenStudy (anonymous):

(5x^2-5x/6x-5)/(x-1/6x-5)

sam (.sam.):

\[\frac{5x^2-5x}{6x-5} \div \frac{x-1}{6x-5}\] Flip the sign will flip the fraction \[\frac{5x^2-5x}{6x-5} \times \frac{6x-5}{x-1}\] Cancel and multiply \[\frac{5x^2-5x}{\cancel{6x-5}} \times \frac{\cancel{6x-5}}{x-1}\] \[(5x^2-5x) \times \frac{1}{(x-1)}\]